# Planning-Driven Programming: A Large Language Model Programming Workflow

### v1

@@ Abstract
The strong performance of large language models (LLMs) on natural language processing tasks raises extensive discussion on their application to code generation. Recent work suggests multiple sampling approaches to improve initial code generation accuracy or program repair approaches to refine the code. However, these methods suffer from LLMs’ inefficiencies and limited reasoning capacity. In this work, we propose an LLM programming workflow (LPW) designed to improve both initial code generation and subsequent refinements within a structured two-phase workflow. Specifically, in the solution generation phase, the LLM first outlines a solution plan that decomposes the problem into manageable sub-problems and then verifies the generated solution plan through visible test cases. Subsequently, in the code implementation phase, the LLM initially drafts a code according to the solution plan and its verification. If the generated code fails the visible tests, the plan verification serves as the intended natural language solution to consistently inform the refinement process for correcting bugs. We further introduce SLPW, a sampling variant of LPW, which initially generates multiple solution plans and plan verifications, produces a program for each plan and its verification, and refines each program as necessary until one successfully passes the visible tests. Compared to the state-of-the-art methods across various existing LLMs, our experimental results show that LPW significantly improves the Pass@1 accuracy by up to 16.4% on well-established text-to-code generation benchmarks, especially with a notable improvement of around 10% on challenging benchmarks. Additionally, SLPW demonstrates up to a 5.6% improvement over LPW and sets new state-of-the-art Pass@1 accuracy on various benchmarks, e.g., 98.2% on HumanEval, 84.8% on MBPP, 64.0% on APPS, and 35.3% on CodeContest, using the advanced LLM GPT-4o as the backbone.
@@ 1 Introduction
Code generation, also known as program synthesis, studies the automatic construction of a program that satisfies a specified high-level input requirement (Gulwani et al., 2017). Recently, large language models (LLMs) pre-trained on extensive code-related datasets (Brown et al., 2020; Meta, 2024; Li et al., 2023; Roziere et al., 2023; Achiam et al., 2023; Muennighoff et al., 2023) have shown success in code-related tasks, such as code generation from natural language descriptions, also named as text-to-code generation (Chen et al., 2021; Austin et al., 2021; Li et al., 2022), code translation (Pan et al., 2024; Yang et al., 2024), and code completion (Izadi et al., 2024). However, LLM-based code generation remains challenging due to stringent lexical, grammatical, and semantic constraints (Scholak et al., 2021). To overcome these challenges, multiple initial programs are generated (Chen et al., 2021; Chowdhery et al., 2023), followed by different best-program selection strategies to improve code generation performance over LLMs (Li et al., 2022; Chen et al., 2023a; Zhang et al., 2023; Ni et al., 2023).
@@
Code generation substantially benefits from the empirical insights of human programmers. In practice, human programmers develop high-quality code by consistently identifying and rectifying errors through the analysis of test case executions, rather than a single effort (Huang et al., 2023c; Chen et al., 2023b). Different studies have refined programs based on execution results and LLM-generated information such as code and error explanation (Tang et al., 2023; Shinn et al., 2023; Madaan et al., 2023). Recent work further optimizes refinement (debugging) methods by performing rubber duck debugging processes (Chen et al., 2023b) and leveraging control flow graph information to assist LLMs in locating bugs (Zhong et al., 2024).
Software development models such as Waterfall and Scrum underscore the importance of communication among various development roles in the production of high-quality software (Davis, 2012; Schwaber, 2004; Andrei et al., 2019). Motivated by this principle, several studies (Lin et al., 2024; Qian et al., 2024; Dong et al., 2023b) have employed LLM instances as customized agents, assigning them diverse development roles and facilitating their collaboration. Multi-agent collaborative code generation emphasizes the distinct workload for each LLM agent, e.g., requirement analyst, architect, programmer, and tester (Lin et al., 2024). Additionally, various communication strategies have been proposed to ensure program quality. For example, Hong et al. (2024) introduced a communication protocol to ensure efficient interactions. Qian et al. (2024) described a communicative dehallucination mechanism to encourage high-quality communication among LLM agents.
However, all the aforementioned methods have certain weaknesses. Multiple sampling approaches suffer from sampling inefficiency and conflict with human programming strategies. In code refinement, feedback messages often lack precise correction instructions, leading to numerous refinements that deviate from the intended solution. Additionally, refining programs that significantly diverge from the problem description remains an open challenge (Tian & Chen, 2023). In multi-agent collaborative code generation, ineffective feedback mechanisms degrade communication quality. This issue is exacerbated when an excessive number of agents are involved, resulting in increased token consumption (Huang et al., 2023a).
Figure 1: The pipeline of LPW, a large language model programming workflow, where the components highlighted in red are exclusive to LPW. LPW consists of two phases. In the solution generation phase, LPW initially creates a solution plan (block (b)) for a problem (block (a)), along with the plan verification (block (c)) for each visible test. If the plan verification infers the accurate output for each visible test based on the solution plan (block (c)) and no incorrect logic is found in the verification check process (block (d)), LPW uses the generated plan and plan verification to help LLMs draft the initial program (block (e)) at the beginning of the code implementation phase. If the initial program passes all visible tests after execution (block (f)), it is used as the final code (block (l)) and then assessed with hidden tests. Otherwise, the LLM-generated code explanation (block (g)) and error analysis (block (j)) serve as debugging inputs to refine the error program (block (k)). The LLM-generated error analysis involves comparing the execution trace (block (h)) with the plan verification (block (i)) on the failed visible test to identify logic flaws in the code implementation and provide repair suggestions. The refined program is reevaluated on the visible tests to determine the necessity for further debugging iterations.
@@
In this work, we propose LPW, a large language model programming workflow, specifically for text-to-code generation, addressing the aforementioned limitations. LPW emulates Waterfall and Scrum where analysis and design steps are prioritized before code implementation. However, LPW streamlines traditional software development models into two phases: the solution generation phase for plan and plan verification creation, and the code implementation phase for initial code development and subsequent refinements. The pipeline of LPW is depicted in Figure 1. LPW leverages various information, including LLM-generated solution plan (Jiang et al., 2023) (block (b)), LLM-generated code explanation (Chen et al., 2023b) (block (g)), and runtime information from program execution (Zhong et al., 2024) (block (h)) to boost the code generation performance, and efficiently incorporates them into an end-to-end framework. In LPW, aside from runtime information, all other messages are autonomously generated by LLMs using few-shot prompting, without the need for annotated corpora or additional training.
@@
A unique feature of LPW is the incorporation of the plan verification (block (c) in Figure 1) as the natural language intended solution for visible tests to derive the reliable program solution. LPW initially produces a solution plan that decomposes complex programs into several tractable sub-problems (intermediate steps) (Cheng et al., 2023; Zelikman et al., 2023; Jiang et al., 2023). LPW then verifies the solution plan against visible tests to assess its correctness, known as plan verification. For a visible test, the verification includes a text-based, step-by-step analysis to derive the output for each intermediate step and the final output, ensuring that the final output is consistent with the visible test result. Additionally, each inferred intermediate output is reviewed by LLMs (block (d) in Figure 1) to maintain logical consistency throughout the verification.
@@
Different from other approaches that exclude the solution plan entirely from the code generation (Chen et al., 2023b; Zhong et al., 2024), LPW incorporates the LLM-generated plan and its verification in the initial code development to clarify the programming logic. This approach ensures that the initial code closely aligns with the problem description, thus reducing the need for subsequent refinements. The plan verification encompasses comprehensive conditions and logical specifications for solving visible tests, eliminating potential misunderstandings before code generation. This is akin to Test-Driven Development, where human developers validate the intended solution with test cases (Beck, 2022). Furthermore, LPW consistently integrates plan verification in the subsequent refinements. In contrast to previous studies (Chen et al., 2023b; Zhong et al., 2024; Shinn et al., 2023) that query LLMs to infer errors in the generated code when it fails a visible test, LPW prompts LLMs to compare the expected output of each intermediate step for solving the failed visible test, as recorded in the plan verification, against the execution trace on the failed visible test to identify discrepancies and further produce an error analysis (block (j) in Figure 1). This approach is more straightforward and reduces uncertainty. These discrepancies assist LLMs in accurately locating bugs and identifying logic flaws in the code implementation, and generating detailed refinement suggestions, as documented in the error analysis. Then, the error analysis when integrated with the code explanation serves as feedback to refine the code in LPW, surpassing conventional scalar or vector rewards and thereby improving the efficiency and accuracy of the refinement process.
@@
We further explore a sampling variant of LPW named as SLPW. SLPW leverages the Upper Confidence Bound (UCB) algorithm (Auer et al., 2002) to balance the exploration and exploitation in debugging multiple generated code samples for optimizing overall performance. We evaluate LPW and SLPW on four text-to-code generation benchmarks: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and their extended test case variants, HumanEval-ET and MBPP-ET (Dong et al., 2023a). We conduct experiments on the proprietary LLM GPT-3.5 (Achiam et al., 2023), and open-source LLMs, Llama-3 (Meta, 2024) and Phi-3 (Abdin et al., 2024). The Pass@1 accuracy (Chen et al., 2021) is reported. The experiment results demonstrate that LPW and SLPW consistently improve text-to-code generation performance across all benchmarks and LLM backbones. Compared to the state-of-the-art LLM debugger, LDB (Zhong et al., 2024), LPW improves Pass@1 accuracy by around 4% across all benchmarks with the GPT-3.5 backbone and achieves up to 16.4% improvement on MBPP when using Llama-3 as the backbone. SLPW shows an additional 1% improvement over LPW with GPT-3.5 and increases accuracy by up to 5.6% over LPW on MBPP with Phi-3. When tested with the advanced GPT-4o (OpenAI, 2024) backbone, LPW and SLPW maintain their advantages, and SLPW achieves new state-of-the-art performance across all benchmarks. Notably, on two challenging benchmarks, APPS (Hendrycks et al., 2021) and CodeContests (Li et al., 2022), LPW and SLPW improve Pass@1 accuracy by around 10% and 5%, respectively, compared to LDB with the GPT-4o backbone.
We outline the key contributions in this paper as follows:
@@
  • We introduce an end-to-end large language model programming workflow, LPW, that draws inspiration from conventional software development models while streamlining and tailoring them specifically for text-to-code generation. LPW significantly improves the code generation accuracy over the state-of-the-art methods.
@@
  • In LPW, we derive the intended solution for visible tests, represented by the plan verification, through querying LLMs to validate the correctness of the LLM-generated solution plan on visible tests before code implementation. The plan verification clarifies all conditions, flow logic, arithmetic operations, and punctuation specifications required to solve the visible tests for the given problem, thereby increasing the LLMs’ confidence during both the initial program generation and subsequent debugging processes.
  • We investigate SLPW, a sampling variant of LPW, and show that debugging across multiple program samples can further enhance performance and set new state-of-the-art results.
@@
  • We conduct extensive experiments across six text-to-code generation benchmarks to validate the performance of LPW and SLPW with various LLM backbones, provide a comprehensive analysis of their performance and failure cases, and highlight the existing challenges.
@@ 2 Problem Formulation
We follow the problem formulation for text-to-code generation as outlined in Jiang et al. (2023), Chen et al. (2023b), and Zhong et al. (2024). The text-to-code generation problem is formulated as a triple $\mathcal{P}=\langle Q,T_{v},T_{h}\rangle$, where $Q$ represents the problem specifications described in natural language, and $T_{v}$ and $T_{h}$ are sets of visible and hidden tests, each containing input-output pairs $(t^{i},t^{o})\in T=T_{v}\cup T_{h}$. The goal is to leverage the LLM $\mathcal{M}$ to generate a program function $f$, $\mathcal{M}\rightarrow f$, that maps each input $t^{i}$ to its corresponding output $t^{o}$ for all pairs in $T$, i.e., $f(t^{i})=t^{o}$, for $(t^{i},t^{o})\in T$. We note that $T_{h}$ remains hidden during both solution generation and code implementation phases and only becomes visible if the generated $f$ passes $T_{v}$. In LPW, for all components shown in Figure 1, the problem description $Q$ is, by default, concatenated with task-specific prompts to produce the desired response from LLMs.
@@ 3 Workflow Structure
Solution Generation. Figure 2 displays the overall workflow of the solution generation phase in LPW (part (a)), with an example programming problem for illustration (part (b)). LPW leverages the self-planning approach introduced by Jiang et al. (2023) to abstract and decompose the problem description $Q$ into a strategic and adaptable plan $\Pi$ at the start of the solution generation phase. For a problem in HumanEval described by block (1) in Figure 2, its example solution plan is illustrated at block (3). However, the LLM-generated plan $\Pi$ may occasionally be incorrect, misguiding subsequent program generation. To avoid this, LPW queries the LLM to verify $\Pi$ against all visible tests $T_{v}$. The LLM-responded plan verification $\mathcal{A}(\Pi,T_{v})$ delivers a step-by-step analysis, including all intermediate results and final derived outputs for all visible tests $T_{v}$ based on $\Pi$. For each $t_{v}\in T_{v}$, its verification $\mathcal{A}(\Pi,\{t_{v}\})$ compares the derived output $t_{v}^{o^{\prime}}$ with the ground-truth output $t_{v}^{o}$ to assess the correctness of $\Pi$, as outlined at block 4 in Figure 2. If $\Pi$ is successfully verified on all visible tests, where in $\mathcal{A}(\Pi,T_{v})$, $t_{v}^{o^{\prime}}=t_{v}^{o},\forall t_{v}\in T_{v}$, then the plan verification $\mathcal{A}(\Pi,T_{v})$ is reviewed by the LLM again to ensure the accuracy of all intermediate results, since each intermediate step result is used in locating bugs and providing refinement suggestions when compared with the code runtime information on the failed visible test. If all intermediate outputs in $\mathcal{A}(\Pi,T_{v})$ are validated as correct by the LLM as shown at block 5 in Figure 2, $\mathcal{A}(\Pi,T_{v})$ is treated as the intended solution for $T_{v}$. The plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$ serve as the output of the solution generation phase, guiding code development and refinements in the code implementation phase.
@@
Figure 3: (a): An illustrated workflow of the code implementation phase in LPW. (b): Example message fragments extending from Figure 2 and corresponding to each workflow component. See Section 6 for detailed messages.
@@
Code Implementation. Figure 3 shows the overall workflow of the code implementation phase in LPW (part (a)), using the same problem from Figure 2 as an illustration (part (b)). LPW develops an initial program $f$ by prompting the LLM with the problem description $Q$ (block (1) in Figure 2), along with plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$ from the solution generation phrase. Subsequently, LPW queries the LLM to add print statements for each line in $f$, resulting in $f_{p}$, and then executes $f_{p}$ on visible tests $T_{v}$. If $f_{p}$ successfully solves $T_{v}$, LPW validates it on the hidden tests $T_{h}$ to report Pass@1 accuracy. Otherwise, LPW collects the runtime information on the first failed visible test $\bar{t_{v}}$, indicating that the implementation in $f$ deviates from the specifications in $\mathcal{A}(\Pi,\{\bar{t_{v}}\})$. Blocks 1-3 in part (b) of Figure 3 depict an initial program $f$ (block (1)) that fails on a visible test $\bar{t_{v}}$ (block (2)) and its execution trace (block (3)) on $\bar{t_{v}}$ after adding print statements. We omit $f_{p}$ from Figure 3 to keep the discussion concise. LPW instructs the LLM to conduct an error analysis by identifying inconsistencies between the intermediate outputs recorded in the execution trace of $\bar{t_{v}}$ and the expected outputs documented in the verification $\mathcal{A}(\Pi,\{\bar{t_{v}}\})$, analyzing the causes, and offering refinement suggestions (block (4)). Subsequently the error analysis and code explanation for $f$ generated by the LLM (block (5)) are concatenated as the prompt to generate the refined program $f^{\prime}$ (block (6)). The code explanation helps the LLM align the text-based error analysis with the code implementation. LPW replaces $f$ with the refined program $f^{\prime}$ and revalidates the updated $f$ against the visible tests $T_{v}$ to assess the necessity for further refinements.
@@
Iterative Updates. LPW includes two update steps in the solution generation phase to enable self-correction as indicated by the red arrows in Figure 2: 1) when the plan verification inferred ultimate output differs from the ground-truth output for a visible test, where $t_{v}^{o^{\prime}}\neq t_{v}^{o},\exists t_{v}\in T_{v}$ in $\mathcal{A}(\Pi,T_{v})$, a revised solution plan $\Pi^{\prime}$ is included in the LLM response to substitute the original plan; 2) when the LLM detects any incorrect intermediate values in $\mathcal{A}(\Pi,T_{v})$ e.g., contextual inconsistencies, mathematical miscalculations, or logical reasoning errors, LPW prompts the LLM to regenerate the plan verification. These update methods ensure that the solution plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$ achieve the necessary precision, as well-formed $\Pi$ and $\mathcal{A}(\Pi,T_{v})$ are essential for the subsequent code generation accuracy (Jiang et al., 2023). In the code implementation phase, the code refinement process acts as an update mechanism, replacing the program $f$ with the refined program $f^{\prime}$ when $f$ fails the visible test $T_{v}$ as highlighted by the red arrow in Figure 3. Overall, for a problem $\mathcal{P}$, LPW iteratively revises the generated plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$, in the solution generation phase, until $\mathcal{A}(\Pi,T_{v})$ infers the correct outputs for all visible tests $T_{v}$ and no error intermediate outputs are present in $\mathcal{A}(\Pi,T_{v})$. Otherwise, LPW reports a failure for $\mathcal{P}$ when reaching the maximum iterations. Similarly, in the code implementation phase, LPW iteratively refines the generated program $f$ if bugs exist. This process continues until a refined $f$ successfully solves $T_{v}$, followed by Pass@1 accuracy calculation on hidden tests $T_{h}$, or LPW reports a failure for $\mathcal{P}$ upon reaching the maximum iteration limit.
4 LPW with Sampling
Text-to-code generation benefits from both multiple sampling and debugging. These two approaches have evolved orthogonally. We propose a sampling variant of LPW, referred as SLPW. SLPW follows the same workflow and update mechanism as LPW but incorporates multiple plan samples $\{\Pi_{1},\ldots\Pi_{k}\}$ and program samples $\{f_{1},\ldots f_{q}\}$. SLPW generates $k$ plan samples at the beginning of the solution generation phase. For each iteration, SLPW leverages the UCB algorithm to competitively select a plan $\Pi$ with the highest upper confidence interval. Then, SLPW performs the same verification process as LPW for $\Pi$. When SLPW verifies $\Pi$ over each visible test and the verification fails on a visible test, it uses the number of visible tests where the plan verification derives an accurate final output as a reward to update the confidence interval of $\Pi$, and $\Pi$ is replaced with the revised plan $\Pi^{\prime}$. Alternatively, when SLPW checks the correctness of intermediate outputs in the plan verification for each visible test and encounters erroneous values, it uses the number of visible tests where the plan verification contains correct intermediate outputs as a reward to update the confidence interval of $\Pi$. SLPW outputs the first $q$, where $q\leq k$, solution plans along with their verifications for the subsequent code implementation phase when solution plans and their verifications are confirmed as correct within the iteration threshold. Otherwise, SLPW provides $[0,q)$ solution plans and their verifications as the output after reaching the iteration threshold. In the code implementation phase, SLPW initially generates a program for each plan and its verification. If no initial program solves $T_{v}$, SLPW applies the UCB algorithm to optimize refinements across multiple programs. It selects a program $f$, refines it, and updates the confidence interval of $f$ based on the number of passed visible tests, until a refined program addresses $T_{v}$ and reports the Pass@1 accuracy on $T_{h}$. Otherwise, the process terminates with a failure upon reaching the iteration threshold. The algorithm details are available in the Appendix A.1.

### v2

@@ Abstract
The strong performance of large language models (LLMs) raises extensive discussion on their application to code generation. Recent research suggests continuous program refinements through visible tests to improve code generation accuracy in LLMs. However, these methods suffer from LLMs’ inefficiency and limited reasoning capacity. In this work, we propose an LLM programming workflow (LPW) designed to improve both initial code generation and subsequent refinements within a structured two-phase workflow. Specifically, the solution generation phase formulates a solution plan, which is then verified through visible tests to specify the intended natural language solution. Subsequently, the code implementation phase drafts an initial code according to the solution plan and its verification. If the generated code fails the visible tests, the plan verification serves as the intended solution to consistently inform the refinement process for correcting bugs. Compared to state-of-the-art methods across various existing LLMs, LPW significantly improves the Pass@1 accuracy by up to 16.4% on well-established text-to-code generation benchmarks. LPW also sets new state-of-the-art Pass@1 accuracy, achieving 98.2% on HumanEval, 84.8% on MBPP, 59.3% on LiveCode, 62.6% on APPS, and 34.7% on CodeContest, using GPT-4o as the backbone.
Planning-Driven Programming: A Large Language Model Programming Workflow
Chao Lei, Yanchuan Chang, Nir Lipovetzky, Krista A. Ehinger School of Computing and Information Systems, The University of Melbourne, Australia {clei1, yanchuanc}@student.unimelb.edu.au, {nir.lipovetzky, kris.ehinger}@unimelb.edu.au
Figure 1: The pipeline of LPW, where the components highlighted in red are exclusive to LPW.
@@ 1 Introduction
Code generation, also known as program synthesis, studies the automatic construction of a program that satisfies a specified high-level input requirement (Gulwani et al., 2017). Recently, large language models (LLMs) pre-trained on extensive code-related datasets (Brown et al., 2020; Meta, 2024; Li et al., 2023; Roziere et al., 2023; Achiam et al., 2023; Muennighoff et al., 2023) have shown success in code-related tasks, such as code generation from natural language descriptions, also named as text-to-code generation (Chen et al., 2021; Austin et al., 2021; Li et al., 2022), code translation (Pan et al., 2024; Yang et al., 2024), and code completion (Izadi et al., 2024). However, LLM-based code generation remains challenging due to stringent lexical, grammatical, and semantic constraints (Scholak et al., 2021).
@@
Code generation substantially benefits from the empirical insights of human programmers. In practice, human programmers develop high-quality code by consistently rectifying errors through the analysis of test case executions (Huang et al., 2023c; Chen et al., 2023b). Different studies have refined programs based on execution results and LLM-generated information such as code and error explanation (Tang et al., 2023; Shinn et al., 2023; Madaan et al., 2023). Recent work further optimizes refinement (debugging) by performing rubber duck debugging processes (Chen et al., 2023b) and leveraging control flow graph information to assist LLMs in locating bugs (Zhong et al., 2024). However, the absence of precise correction instructions in feedback messages results in numerous refinements that deviate from the intended solution. Additionally, refining programs that significantly diverge from the problem description remains an open challenge (Tian and Chen, 2023).
To replicate each phase of program development, several studies (Lin et al., 2024; Qian et al., 2024; Dong et al., 2023b) have employed LLM instances as customized agents, assigning them diverse roles and facilitating their collaboration. Recent work incorporates extra visible tests (Huang et al., 2023a) and solution plans (Islam et al., 2024) generated by designated agents to improve the code refinements in multi-agent collaborations. However, the absence of methodologies for generating reliable visible tests and plans in these studies undermines their credibility, as incorrect visible tests and plans can lead to erroneous codes. Besides, with an increased number of agents, multi-agent collaborations consume significant token resources for communication (Huang et al., 2023a). The detailed related work is discussed in Appendix A.
@@
In this work, we propose LPW, a large language model programming workflow, specifically for text-to-code generation, addressing the aforementioned limitations. LPW involves two phases for code generation: the solution generation phase for plan creation and plan verification, and the code implementation phase for initial code development and subsequent refinements. The pipeline of LPW is depicted in Figure 1. LPW leverages various information, including LLM-generated solution plan (Jiang et al., 2023) (block (b)), LLM-generated code explanation (Chen et al., 2023b) (block (g)), and runtime information from program execution (Zhong et al., 2024) (block (h)) to boost the code generation performance, and effectively incorporates them into an end-to-end framework. In LPW, aside from runtime information, all other messages are autonomously generated by LLMs using few-shot prompting, without the need for annotated corpora or additional training.
@@
A unique feature of LPW is incorporating plan verification (block (c)) as the natural language intended solution for visible tests. LPW initially produces a solution plan that decomposes a complex problem into several tractable sub-problems (intermediate steps) (block (b)). LPW then verifies the solution plan against visible tests to assess its correctness, known as plan verification. For a visible test, the LLM-generated verification includes a text-based step-by-step analysis to derive the output for each intermediate step and the final output, ensuring that the final output is consistent with the visible test result. Subsequently, each inferred intermediate output is reviewed by LLMs (block (d)) to maintain logical consistency and mitigate hallucination throughout the verification. The plan verification encompasses comprehensive conditions and logical specifications for solving visible tests, eliminating potential misunderstandings before code generation. This is akin to Test-Driven Development, where human developers validate the intended solution with test cases (Beck, 2022).
@@
LPW uses the plan and its verification in the initial code development (block (e)) to ensure that the initial code closely aligns with the problem description. Furthermore, LPW incorporates plan verification in the subsequent refinements. By comparing discrepancies between the expected output of each intermediate step, as recorded in the plan verification, against the execution trace on the failed visible test (block (h)), LPW accurately locates bugs, identifies logic flaws in the code implementation, and further generates detailed refinement suggestions, as documented in the error analysis (block (j)). Then, the error analysis when integrated with the code explanation (block (g)) serves as feedback to refine the code, surpassing conventional scalar rewards or deduced error analysis (Chen et al., 2023b; Zhong et al., 2024; Shinn et al., 2023) and thereby improving refinement efficiency and accuracy.
@@
We first evaluate LPW on four text-to-code generation benchmarks: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and their extended test case variants, HumanEval-ET and MBPP-ET (Dong et al., 2023a). We conduct experiments on the proprietary LLM GPT-3.5 (Achiam et al., 2023), and open-source LLMs, Llama-3 (Meta, 2024) and Phi-3 (Abdin et al., 2024). The Pass@1 accuracy (Chen et al., 2021) is reported. Compared to the state-of-the-art LLM debugger, LDB (Zhong et al., 2024), LPW improves Pass@1 accuracy by 2% to 6.1% across all benchmarks with GPT-3.5 and achieves up to 16.4% improvement on MBPP with Llama-3. When evaluated on additional benchmarks using the advanced GPT-4o (OpenAI, 2024), LPW maintains its advantages, and achieves new state-of-the-art performance across evaluated benchmarks. Notably, on the contamination-free benchmark, LiveCode (Jain et al., 2024), and challenging benchmarks, APPS (Hendrycks et al., 2021) and CodeContests (Li et al., 2022), LPW improves Pass@1 accuracy by around 5%, 10%, and 5%, respectively, compared to LDB. We outline our contributions as follows:
@@
  • We introduce an end-to-end large language model programming workflow, LPW, which significantly improves the code generation accuracy over the state-of-the-art methods.
@@
  • We derive the intended solution for visible tests, represented as the plan verification. The plan verification clarifies logic specifications required to solve the visible tests for the given problem, thereby increasing the LLMs’ confidence during both the initial program generation and subsequent debugging processes.
@@
  • We conduct extensive experiments across seven text-to-code generation benchmarks to validate the performance of LPW with various LLM backbones, provide a comprehensive analysis of their performance, token usage, and failure cases, and highlight the existing challenges.
@@ 2 Problem Formulation
We follow the problem formulation for text-to-code generation as outlined in Jiang et al. (2023), Chen et al. (2023b), and Zhong et al. (2024). The text-to-code generation problem is formulated as a triple $\mathcal{P}=\langle Q,T_{v},T_{h}\rangle$, where $Q$ represents the natural language problem specifications, and $T_{v}$ and $T_{h}$ are sets of visible and hidden tests, each containing input-output pairs $(t^{i},t^{o})\in T=T_{v}\cup T_{h}$. The goal is to leverage the LLM $\mathcal{M}$ to generate a program function $f$, $\mathcal{M}\rightarrow f$, that maps each input $t^{i}$ to its output $t^{o}$ for all pairs in $T$, i.e., $f(t^{i})=t^{o}$, for $(t^{i},t^{o})\in T$. We note that $T_{h}$ remains hidden during both solution generation and code implementation phases and only becomes visible if the generated $f$ passes $T_{v}$. In LPW, for all components shown in Figure 1, the problem description $Q$ is, by default, concatenated with task-specific prompts to produce the desired response from LLMs.
@@ 3 Workflow Structure
Solution Generation. Figure 2 displays the overall workflow of the solution generation phase in LPW (part (a)), with an example programming problem for illustration (part (b)). LPW leverages the self-planning approach introduced by Jiang et al. (2023) to abstract and decompose the problem description $Q$ into a strategic and adaptable plan $\Pi$ at the start of the solution generation phase.
@@
For a problem in HumanEval described by block (1) in Figure 2, its example solution plan is illustrated at block (3). However, the LLM-generated plan $\Pi$ may occasionally be incorrect, misguiding subsequent program generation. To avoid this, LPW queries the LLM to verify $\Pi$ against all visible tests $T_{v}$. The LLM-responded plan verification $\mathcal{A}(\Pi,T_{v})$ delivers a step-by-step analysis, including all intermediate results and final derived outputs for all visible tests $T_{v}$ based on $\Pi$. For each $t_{v}\in T_{v}$, its verification $\mathcal{A}(\Pi,\{t_{v}\})$ compares the derived output $t_{v}^{o^{\prime}}$ with the ground-truth output $t_{v}^{o}$ to assess the correctness of $\Pi$, as outlined at block 4. If $\Pi$ is successfully verified on all visible tests, where in $\mathcal{A}(\Pi,T_{v})$, $t_{v}^{o^{\prime}}=t_{v}^{o},\forall t_{v}\in T_{v}$, then the plan verification $\mathcal{A}(\Pi,T_{v})$ is reviewed by the LLM again to ensure the accuracy of all intermediate results, since each intermediate step result is used in locating bugs and providing refinement suggestions when compared with the code runtime information on the failed visible test. If all intermediate outputs in $\mathcal{A}(\Pi,T_{v})$ are validated as correct by the LLM as shown at block 5, $\mathcal{A}(\Pi,T_{v})$ is treated as the intended solution for $T_{v}$. The plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$ serve as the output of the solution generation phase, guiding code development and refinements in the code implementation phase.
@@
Figure 3: (a): An illustrated workflow of the code implementation phase in LPW. (b): Example message fragments extending from Figure 2 and corresponding to each workflow component. See Section 6 for detailes.
@@
Code Implementation. Figure 3 shows the overall workflow of the code implementation phase in LPW (part (a)), using the same problem from Figure 2 as an illustration (part (b)). LPW develops an initial program $f$ by prompting the LLM with the problem description $Q$ (block (1) in Figure 2), along with plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$ from the solution generation phase. Subsequently, LPW queries the LLM to add print statements for each line in $f$, resulting in $f_{p}$, and then executes $f_{p}$ on visible tests $T_{v}$. If $f_{p}$ successfully solves $T_{v}$, LPW validates it on the hidden tests $T_{h}$ to report Pass@1 accuracy. Otherwise, LPW collects the runtime information on the first failed visible test $\bar{t_{v}}$, indicating that the implementation in $f$ deviates from the specifications in $\mathcal{A}(\Pi,\{\bar{t_{v}}\})$. Blocks 1-3 in part (b) of Figure 3 depict an initial program $f$ (block (1)) that fails on a visible test $\bar{t_{v}}$ (block (2)) and its execution trace (block (3)) on $\bar{t_{v}}$ after adding print statements. We omit $f_{p}$ from Figure 3 to keep the discussion concise. LPW instructs the LLM to conduct an error analysis by identifying inconsistencies between the intermediate outputs recorded in the execution trace of $\bar{t_{v}}$ and the expected intermediate outputs documented in the verification $\mathcal{A}(\Pi,\{\bar{t_{v}}\})$, analyzing causes, and offering refinement suggestions (block (4)). Subsequently, the error analysis and code explanation for $f$ generated by the LLM (block (5)) are concatenated as the prompt to generate the refined program $f^{\prime}$ (block (6)). The code explanation helps the LLM align the text-based error analysis with the code implementation. LPW replaces $f$ with the refined program $f^{\prime}$ and revalidates the updated $f$ against the visible tests $T_{v}$ to assess the need for further refinements.
@@
Iterative Updates. LPW includes two update steps in the solution generation phase to enable self-correction as indicated by the red arrows in Figure 2: 1) when the plan verification inferred final output differs from the ground-truth output for a visible test, where $t_{v}^{o^{\prime}}\neq t_{v}^{o},\exists t_{v}\in T_{v}$ in $\mathcal{A}(\Pi,T_{v})$, a revised solution plan $\Pi^{\prime}$ is included in the LLM response to substitute the original plan; 2) when the LLM detects any incorrect intermediate values in $\mathcal{A}(\Pi,T_{v})$ e.g., contextual inconsistencies, mathematical miscalculations, or logical flaws, LPW prompts the LLM to regenerate the plan verification. These update methods ensure that the solution plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$ maintain the necessary precision, as well-formed $\Pi$ and $\mathcal{A}(\Pi,T_{v})$ are essential for accurate code generation (Jiang et al., 2023). In the code implementation phase, the code refinement process acts as an update mechanism, replacing the program $f$ with the refined program $f^{\prime}$ when $f$ fails the visible test $T_{v}$ as highlighted by the red arrow in Figure 3. Overall, for a problem $\mathcal{P}$, LPW iteratively revises the generated plan $\Pi$ and its verification $\mathcal{A}(\Pi,T_{v})$, in the solution generation phase, until $\mathcal{A}(\Pi,T_{v})$ infers correct outputs for all visible tests $T_{v}$ and no error intermediate outputs are present in $\mathcal{A}(\Pi,T_{v})$. Otherwise, LPW reports a failure for $\mathcal{P}$ when reaching the maximum iterations. Similarly, in the code implementation phase, LPW iteratively refines the generated program $f$ if bugs exist. This process continues until a refined $f$ successfully solves $T_{v}$, followed by Pass@1 accuracy calculation on hidden tests $T_{h}$, or LPW reports a failure for $\mathcal{P}$ upon reaching the maximum iteration limit.